# Quantum Computing in Wireless Communications and Networking: A Tutorial-cum-Survey

### v1

@@ III-B3 Applications of AI-driven QC for Power Allocation
The previous studies on power allocation rely on conventional QC algorithms like QPSO. However, it is worth noting that these conventional algorithms still demand significant computational resources and lead to high overhead, such as computation delays. Recent research has turned towards AI-driven QC solutions for power allocation. These approaches, such as QNN and QDRL, leverage advanced machine learning techniques to mitigate the computational complexities inherent in traditional QC algorithms. By harnessing the AI power, these methods offer promising avenues for more efficient and effective power allocation in wireless networks. For example, [120] proposed a distributed QNN framework for the purpose of optimizing wireless communication and gave a case study in non-orthogonal multiple access system for the framework implementation. The objective was also to maximize the weighted-sum rate under consideration of channel interference, and the proposed QNN algorithm was implemented in quantum circuits, which accelerates running time overhead.
Other works implemented in quantum computers include [121, 122] in the field of power allocation. [121] presented a QNN with parallel training for optimizing wireless resource allocation. The proposal reduced the dimension of training data by sending only the statistical parameters of the dataset instead of the whole dataset. The proposed method was applied to optimize transmit precoding and power allocation in MIMO-NOMA. The analysis showed that the parallel training approach achieved a comparable sum rate with lower complexity compared to conventional methods. It highlighted the potential benefits of QNNs in processing large dimensional solution spaces and addressed the challenges of distributed datasets and training complexity. [122] proposed a layerwise QDRL method for optimizing continuous large space and time series problems using deep layer training. It leveraged the advantages of QC to maximize reward and reduce training loss. The study focused on jointly optimizing UAV trajectory planning, user grouping, and power allocation for higher energy efficiency. The main contributions of the study were the introduction of layerwise quantum embedding, and the presentation of quantum circuits for practical implementation. However, it was assumed that users can perceive perfect CSI in the UAV downlink communication channel, which limits its application.
@@
In conclusion, the integration of QC techniques in power allocation holds promise for optimizing various aspects of wireless networks. Table III provides an overview of the main findings, optimization objectives, and QC algorithms utilized across different wireless networks. Noting that, there are relatively fewer studies employing QC algorithms for power allocation, largely due to the localized nature of power allocation in wireless networks, typically confined to a single-cell system. This observation is reflected in both the choice of QC algorithms and their control manners as depicted in Table IV. The majority of studies utilize QC simulated algorithms that are implemented on traditional computers.
@@ III-C1 Overview
TABLE V: Summary of QC Works on Channel Assignment in Wireless Networks
@@
TABLE VI: Details of QC Algorithms for Channel Assignment
@@ III-C2 Applications of Conventional QC for Channel Assignment
The application of QAOA was discussed to solve the problem of maximum likelihood detection of binary symbols transmitted over MIMO channel [124]. The maximum likelihood detection problem is NP-hard and difficult to solve using classical computers. The study utilized QAOA to encode maximum likelihood detection into a level-p QAOA circuit, which was solved using classical optimizers. The work provided an overview of QAOA and their applications in solving various problems.
@@
D2D communications with $N$ transmitter-receiver pairs in a single-cell system was discussed in [125]. The authors highlighted the challenge of mutual interference between D2D and nearby cellular communication and the need to reduce the interference through resource allocation including subchannel assignment and power allocation. They formulated the problem as a capacitated Max k-Cut and applied QA. Specifically, sub-channels were assigned to disjoint transmitter groupings, and the issue was formulated as a QUBO problem. The algorithm was implemented in a quantum computer. Though the paper considered a simple scenario of subchannel allocation, it provided guidelines to apply optimization techniques of QC to solve similar problems. Applying another QC algorithm QGA for cognitive radio spectrum allocation, illustrated in Fig. 12, in a single-cell network was discussed in [126]. Specifically, the work proposed a mapping process between the channel assignment matrix and the chromosome of the algorithms to decrease the search space. Thought it highlighted the importance of dynamic spectrum access in order to improve spectrum utilization in the validation results, there is no consideration in the problem formulation and analysis.
@@
The preceding works explored QC algorithms for channel assignment in local wireless networks. Their application in mobile cellular networks was also conducted in [127]. It presented a case study on the root sequence index assignment problem in LTE/NR physical random access channel configuration. They formulated the problem as a QUBO problem and solved it using a cloud-based QC platform. The results showed that QA solver can successfully assign conflict-free root sequence index, but comparison with classical heuristics revealed that some classic algorithms were more effective in terms of solution quality and computation time. The reason lies in the semi-quantum algorithm and the type of quantum computer, which provides reference for the correct use of QC techniques.
QC in relaying networks:
@@
The problem of channel assignment in relaying networks is typically formulated as conflict graphs that were resolved by QA in [128] and [128]. A real-world application of QA, specifically the scheduling of air-links for maximum throughput while avoiding interference near network nodes was studied in [128]. It introduced a novel gap expansion technique that benefits QA more than simulated annealing, leading to improved speedup and network queue occupancy. The proposed QA was implemented in quantum computers. However, it is assumed that overall network information was given in the scheduling algorithm. [129] explored network performance in the same scenario with [128]. Weighted maximum independent set was utilized to model the K-hop interference. It made a comparison between simulated annealing and QA in solving the scheduling problem. The results showed that the QA solution outperformed simulated annealing when using the gap expansion technique.
@@ III-C3 Applications of AI-driven QC for Channel Assignment
A framework of QNNs for resource allocation [130] was designed in wireless communications, such as 5G/6G, to reduce time complexity while maintaining performance, and a reinforcement-learning-inspired QNN was proposed to further improve performance. Specifically, they presented the quantum circuit design of the QNN to ensure practical implementation in NISQ computers. The QNN was applied to group users in NOMA as a case study. The performance on channel assignment outperformed other solutions, but at the expense of a large number of qubits. RL based QC algorithms, quantum multi-agent AC networks, were performed in [131] to construct a robust mobile access system with multiple UAVs. The algorithm utilized MARL to enable UAVs to learn collectively and optimize their actions within a shared environment. The study addressed scalability challenges by implementing a quantum centralized critic. One of its advantages lied in the design of a realistic environment with noise injection, the evaluation of QMARL and noise reflection on training policies.
One of 6G scenarios, Metaverse, was discussed in [132]. The problem lied in insufficient sampling in model training for connected and autonomous vehicles (CAVs). To efficiently achieve Metaverse, channel resources among RSUs and vehicles were significant for network performance, in which eNodeBs allocated channel resources to RSUs and RSUs allocated them to vehicles. It formulated the channel resource assignment as a discrete Markov Decision Process (MDP) and adopted QRL to obtain the optimal policy. However, the study assumed that vehicle information was given without considering their dynamic, such as moving out communication ranges of RSUs.
@@ III-C4 Applications of QC Algorithms in IRS
Intelligent Reflective Surfaces (IRSs) [138] represent a revolutionary technology in 6G communications that leverages passive elements to manipulate electromagnetic waves. These surfaces are composed of programmable materials capable of adjusting the amplitude, phase, and polarization of incident electromagnetic signals as illustrated in Fig. 13a. By intelligently controlling these properties, IRSs can effectively shape and redirect wireless signals to enhance communication performance.
@@
Figure 13: Different IRS systems
@@
[135] proposed an approach to controlling the propagation path characteristics from the transmitter to the receiver by appropriately according the reflection phase of IRS elements in a communication cell consisting of one BS and multiple users under assistance of multiple IRSs as illustrated in Fig. 13b. The core problem was to schedule the assignment of IRSs to a given users for a unit time. The authors first formulated the scheduling problem as a QUBO problem and then developed an algorithm based on QA to solve the issue. The solution was validated in the quantum computers with the platform D-Wave. There are other works [133] and [134] on the same scenario as [135]. Taking [133] as an example, it compared its tensor contraction and regression based QA with PSO and factorization machine based QA. However, there was no explanation for its advantage over others.
@@
[135] and [133] considered the scheduling problem with only one BS in a communication cell. To improve system scalability, [136] tackled the IRS sub-channel assignment with multiple BSs as illustrated in Fig. 13c. The network structure was different from the scenario in which there was only one BS, when multiple BSs were introduced into the network. In this work, IRSs and users were coordinated by a BS and a server (such as a cellular station), but were not necessarily in the same coverage cell of the BS. When there were multiple BSs in different cells under assistance of multiple IRSs, the channel scheduling problem was formulated as a combinatorial optimization, which was also in the area of QUBO. And a QA based solution was proposed to solve the problem. However, it was a centralized scheduling method for multiple IRSs. In addition, there was no detailed discussion on QA implementation though the simulations validated its effectiveness.
@@
[135] and [133] considered network performance in a communication cell with multiple users, an IRS and one BS. The network was easily extended to a communication cell with multiple users, multiple IRSs, and multiple BSs [136] in different cells. To reduce the network overhead such as hardware costs, it obtained the similar network performance with [136] by sharing an ISR of switching phase shift as illustrated in Fig. 13d, which was studied in [137]. Specifically, when a particular phase shift was set at a given time, the corresponding time resources and BSs were assigned to UEs existing in the reflection direction of the radio waves transmitted from each BS, which enabled efficient communication. In the case of multiple operators, each operator performed such radio resource-allocation control in parallel, enabling communications with the IRS without coordination required between the operators. The process was formulated as QUBO and solved by classical QC algorithm. However, it can not achieve the optimal performance since they worked independently. And thus, it is still viewed as a centralized approach. Other works related to IRS scheduling in the area of channel assignment includes [134] that considered both power allocation and communication security. Its scenario was similar with [135] and [133] in a communication cell. However, the solution for channel assignment of scheduling phase shift was independent with other optimization variables. The quantum-inspired bald eagle search algorithm was developed to solve the problem, which was not orthodox QC.
@@
In summary, QC-enabled algorithms offer a promising solution to enhance performance in channel assignment, as the overview of these studies presented in Table V and QC details in Table VI.
@@ III-D1 Overview
User association in wireless networks refers to the process of determining which access point, BS, or sensor sink that a user device connects to for communication. This decision is crucial for optimizing network performance, ensuring efficient resource utilization and service quality [139]. The primary goal of user association is to provide seamless connectivity while maximizing network capacity and throughput. Various factors influence user association decision, including signal strength, available bandwidth, network load, user mobility, and QoS requirements. Note that network dynamics, particularly, sink (such as UAVs) mobility, are one of key features in user association, which was explored in numerous studies [140], such as in mobile sensor networks and UAV networks. Solutions encounter computational constraints like with a substantial number of devices and dynamic environment. The challenge can be addressed through the utilization of QC techniques. Thus, in this section, we primarily categorize our discussion according to two key criteria: QC algorithm category and mobility of BSs or sinks.
@@ III-D2 Applications of Conventional QC for User Association
TABLE VII: Summary of QC Works on User Association in Wireless Networks
@@
The study examineda basic wireless sensor network [141], consisting of sensors and a BS. It introduced a clustering method based on QEA to extend the network lifespan. However, it does not account for network dynamics, such as traffic patterns and device mobility. In [142], a comparable scenario to [141] was explored, with the introduction of the quantum-inspired gravitational search to enhance energy efficiency and prolong network lifespan. The study considered communication constraints and application-specific criteria, encompassing parameters such as mean relative deviation, spatial density error, and sensor-out-of-range error. Nonetheless, the paper lacked an in-depth analysis of the optimality of the proposed solution. Meanwhile, [143] presented an alternative quantum-inspired approach to the issue addressed in [141] and [142]. Named the quantum-inspired gravitational search algorithm, this solution aimed to determine the operational states of sensors while preserving network connectivity. It introduced a weighted fitness function, integrating connectivity, energy, and application-specific parameters. However, accurately determining the appropriate weights for these factors posed a challenge.
Several additional quantum-inspired approaches have been proposed in the literature, including works such as [144, 145, 146, 147, 148, 149]. These contributions introduced techniques such as quantum elite gray wolf optimization, quantum clone whale optimization, quantum artificial bee colony, QPSO, and QGA, all aimed at addressing the same problem outlined in [141]. Additionally, other studies explored link connectivity within QC algorithms. For example, [150] investigated online social networks like Facebook, Weibo, and Twitter. The goal was to predict link connectivity, facilitating recommendations for registered users to connect with others.
@@
The preceding examination of user association in the absence of accounting for mobility and network dynamics has been extensively explored in existing literature. Subsequent studies consider more complex scenarios involving UAV networks and wireless sensor networks featuring mobile sinks. For example, [151] focused on optimizing the path planning of a mobile sink to enhance energy efficiency. In this scenario, the mobile sink started its path from a designated location, traversed through the communication range of each sensor along the predetermined path, conducted a round of data collection, and subsequently returned to its initial starting point. They employed mathematical transformations to align the problem characteristics with those of a traveling salesman (TSP) problem. An optimal TSP path was obtained by using quantum tunneling effect and quantum circuit to achieve parallelism.
Regarding UAV networks, [152] considered a scenario with multiple UAVs but did not address data relay among UAVs. The research designed an objective function incorporating parameters like average distance, distance to neighboring UAVs, and the number of UAV neighbors to select cluster heads. While the proposal accounted for UAV mobility, it lacked mobility analysis within the model. Additionally, the paper explored UAV clustering application in image classification, but the model lacked a direct correlation between UAV clustering and image classification. In contrast, [153] tackled the challenge of UAV coalition formation for providing computation resources in a large-scale, heterogeneous UAV network. The study operated in an unknown and dynamic environment, where UAVs must dynamically acquire environmental knowledge. Leader UAVs assessed follower UAVs’ contributions in resource availability after completing missions, sharing this information during coalition formation. QGA guided the inclusion of idle UAVs into coalitions, considering UAV mobility characterized by randomness and lacking correlation with ground points of interest (PoI). However, the study lacked an analysis of PoI impact on system performance, differing significantly from edge computing research, particularly in task offloading investigations [156].

### v3

@@ III-B3 Applications of AI-driven QC for Power Allocation
The previous studies on power allocation rely on conventional QC algorithms like QPSO. However, it is worth noting that these conventional algorithms still demand significant computational resources and lead to high overhead, such as computation delays. Recent research has turned towards AI-driven QC solutions for power allocation. These approaches, such as QNN and QDRL, leverage advanced ML techniques to mitigate the computational complexities inherent in traditional QC algorithms. By harnessing the AI power, these methods offer promising avenues for more efficient and effective power allocation in wireless networks. For example, [99] proposes a distributed QNN framework for the purpose of optimizing wireless communication and gives a case study in non-orthogonal multiple access system for the framework implementation. The objective is also to maximize the weighted-sum rate under consideration of channel interference, and the proposed QNN algorithm is implemented in quantum circuits, which accelerates running time.
Other works implemented in quantum computers include [100, 101] in the field of power allocation. [100] presents a QNN with parallel training for optimizing wireless resource allocation. The proposal reduces the dimension of training data by sending only the statistical parameters of the dataset instead of the whole dataset. The proposed method is applied to optimize transmit precoding and power allocation in MIMO-NOMA. The analysis shows that the parallel training approach achieves a comparable sum rate with lower complexity compared to conventional methods. It highlights the potential benefits of QNNs in processing large dimensional solution spaces and in addressing the challenges of distributed datasets and training complexity. Specifically, similar to neurons in deep learning, the proposed QNNs for parallel training apply qubits as the basic units of computation. Each qubit, associated with a parameterized gate, is measured at the end of the QNN inference. The measurement yields the result for each individual qubit, and the corresponding loss gradient for the $i$-th qubit is calculated at the same time. This allows all weight parameters to be updated simultaneously, which significantly reduces the overall complexity of the for parallel training process. There are three modules in the QNN operations: encoding, QNN inference, and decoding. The input encoding is performed to map an arbitrary classical value to its quantum state, in which encoding is implemented by a quantum-based operation. The QNN inference is similar to the forward propagation in deep learning. Afterwards, each quantum measurement gives result to either 1 or 0. The result of quantum measurements is then decoded as QNN output by means of assigning quantum states.
[101] proposes a layerwise QDRL method for optimizing continuous large space and time series problems using deep layer training. It leverages the advantages of QC to maximize reward and reduce training loss. The study focuses on jointly optimizing UAV trajectory planning, user grouping, and power allocation for energy efficiency. The main contributions of the study are the introduction of layerwise quantum embedding, and the presentation of quantum circuits for practical implementation. In quantum embedding, the encoding operation typically consists of multiple layers, with quantum measurements performed only after the final layer to obtain classical values. In this study, it adopts a layer-wise approach to training, where quantum measurements are conducted at the end of each layer. The classical values derived from these measurements are then used to compute the local loss for that layer. This localized loss is subsequently used to update the training parameters for the next layer, allowing for more granular optimization during the training process. However, the solution is assumed that users can perceive perfect CSI in the UAV downlink communication channel, which limits its application.
@@
In conclusion, the integration of QC techniques in power allocation holds promise for optimizing various aspects of wireless networks. Table IV provides an overview of the main findings, optimization objectives, and QC algorithms utilized across different wireless networks. Noting that, there are relatively fewer studies employing QC algorithms for power allocation, largely due to the localized nature of power allocation in wireless networks, typically confined to a single-cell system. This observation is reflected in both the choice of QC algorithms and their control manners as depicted in Table V. Note that the algorithms can be classified into orthodox QC algorithms and simulated ones. Orthodox QC algorithms refer to those that are directly implemented on quantum computers, leveraging quantum hardware for execution. The majority of studies utilize QC simulated algorithms that are implemented on traditional computers.
@@ III-C1 Overview
TABLE VI: Summary of QC Works on Channel Assignment in Wireless Networks
@@
TABLE VII: Details of QC Algorithms for Channel Assignment
@@ III-C2 Applications of Conventional QC for Channel Assignment
The application of QAOA is discussed to solve the problem of maximum likelihood detection of binary symbols transmitted over MIMO channels [103]. The maximum likelihood detection problem is NP-hard and difficult to solve by classical computers. The study utilizes QAOA to encode maximum likelihood detection into a level-p QAOA circuit, which is solved by classical optimizers. The work provides an overview of QAOA and their applications in solving various problems.
@@
Device to device (D2D) communications with $N$ transmitter-receiver pairs in a single-cell system are discussed in [104]. The authors highlight the challenge of mutual interference between D2D and nearby cellular communication and the need to reduce the interference through resource allocation including sub-channel assignment and power allocation. They formulate the problem as a capacitated max k-cut and applied quantum annealing (QA). Specifically, sub-channels are assigned to disjoint transmitter groupings, and the issue is formulated as a QUBO problem. The algorithm is implemented in a quantum computer. Though the paper considers a simple scenario of sub-channel allocation, it provides guidelines to apply optimization techniques of QC to solve similar problems. Applying another QC algorithm QGA for cognitive radio spectrum allocation, illustrated in Fig. 12, in a single-cell network is discussed in [105]. Specifically, the work proposes a mapping process between the channel assignment matrix and the chromosome of the algorithms to decrease the search space. Thought it highlights the importance of dynamic spectrum access in order to improve spectrum utilization in the validation results, there is no consideration in the problem formulation and analysis.
@@
The preceding works explore QC algorithms for channel assignment in local wireless networks. Their application in mobile cellular networks is also conducted in [106]. It presents a case study on the root sequence index assignment problem in LTE/NR physical random access channel configuration. They formulate the problem as a QUBO problem and solved it using a cloud-based QC platform. The results show that QA solver can successfully assign conflict-free root sequence index, but comparison with classical heuristics reveals that some classic algorithms are more effective in terms of solution quality and computation time. The reason lies in the semi-quantum algorithm and the type of quantum computer, which provides reference for the correct use of QC techniques.
QC in relay networks:
@@
The problem of channel assignment in relay networks is typically formulated as conflict graphs that are resolved by QA in [107, 108]. A real-world application of QA, specifically the scheduling of air-links for maximum throughput while avoiding interference near network nodes is studied in [107]. It introduces a novel gap expansion technique that benefits QA more than simulated annealing, leading to improved speedup and network queue occupancy. The proposed QA is implemented in quantum computers. However, it is assumed that overall network information is given in the scheduling algorithm. [108] explored network performance in the same scenario with [107]. Weighted maximum independent set is utilized to model the K-hop interference. It makes a comparison between simulated annealing and QA in solving the scheduling problem. The results show that the QA solution outperforms simulated annealing when using the gap expansion technique.
@@ III-C3 Applications of AI-driven QC for Channel Assignment
A framework of QNNs for resource allocation [109] is designed in wireless communications, such as 5G/6G, to reduce time complexity while maintaining performance, and a reinforcement-learning-inspired QNN is proposed to further improve performance. Specifically, they present the quantum circuit design of the QNN to ensure practical implementation in NISQ computers. The proposed QNN architecture comprises three key components: an encoding stage, a multi-layer network, and a decoding stage. The encoding phase transforms the original input and weight values into quantum superposition states. The main operation of the QNN resembles that of a fully connected classical neural network, where the CCX(·) and CZ(·) gates are utilized as activation functions and inter-neuron connections, respectively. The decoding stage is responsible for extracting real-valued outputs from the QNN. Since the loss function is computed using real numbers, ratio decoding is applied to map qubit states to real-number outputs. The QNN is applied to group users in NOMA as a case study. The performance on channel assignment outperforms other solutions, but at the expense of a large number of qubits.
RL based QC algorithms, quantum multi-agent AC networks, are performed in [110] to construct a robust mobile access system with multiple UAVs. The proposed QDRL framework consists of a centralized quantum critic and multiple quantum actor networks. The $m$-th quantum actor selects the action with the highest probability among the available actions, based on its state and observation information. Meanwhile, the centralized critic evaluates the current state using a state-value function. The actor and critic networks are updated using the temporal difference actor-critic model, guided by the Bellman optimality equations. Due to the unique QC properties, such as qubit rotation via unitary operations, multi-qubit entanglement, and complex conjugates in the Q and V functions, the traditional policy gradient derivative can not be computed directly. This is because the quantum state remains unknown prior to measurement. To address this challenge, the parameter-shift rule is applied to the Bellman optimality equations. This approach enables the calculation of gradients by linking classical derivatives with QC, thereby integrating classical RL with QC. The algorithm utilizes MARL to enable UAVs to learn collectively and optimize their actions within a shared environment. The study addresses scalability challenges by implementing a quantum centralized critic. One of its advantages lies in the design of a realistic environment with noise injection, the evaluation of quantum multiple agent reinforcement learning (QMARL) and noise reflection on training policies.
One of 6G scenarios, Metaverse, is discussed in [111]. The problem lies in insufficient sampling in model training for connected and autonomous vehicles. To efficiently achieve Metaverse, channel resources among road-side units (RSUs) and vehicles are significant for network performance, in which eNodeBs allocate channel resources to RSUs and RSUs allocate them to vehicles. It formulates the channel resource assignment as a discrete Markov decision process (MDP) and adopts QRL to obtain the optimal policy. However, the study assumes that vehicle information is given without considering their dynamic, such as moving out communication ranges of RSUs.
@@ III-C4 Applications of QC Algorithms in IRS
Intelligent reflective surfaces (IRSs) [117] represent a revolutionary technology in 6G communications that leverages passive elements to manipulate electromagnetic waves. These surfaces are composed of programmable materials capable of adjusting the amplitude, phase, and polarization of incident electromagnetic signals as illustrated in Fig. 13a. By intelligently controlling these properties, IRSs can effectively shape and redirect wireless signals to enhance communication performance. IRS systems fundamentally enhance wireless communication by reflecting and manipulating signals, which makes them totally different in technique from channel or spectrum assignment. This independent treatment allows us to highlight its unique role in optimizing communication, which is distinct from traditional techniques in spectrum assignment.
@@
Figure 13: Different IRS systems.
@@
[114] proposes an approach to controlling the propagation path from the transmitter to the receiver by appropriately according the reflection phase of IRS elements in a communication cell consisting of one BS and multiple users under assistance of multiple IRSs as illustrated in Fig. 13b. The core problem is to schedule the assignment of IRSs to a given users for a unit time. The authors first formulate the scheduling problem as a QUBO problem and then develop an algorithm based on QA to solve the issue. The solution is validated in quantum computers with the platform D-Wave. There are other works [112, 113] for the same scenario as [114]. Taking [112] as an example, it compares its tensor contraction and regression with PSO and factorization machine based QA. However, there is no explanation for its advantage over others.
@@
[114] and [112] consider the scheduling problem with only one BS in a communication cell. To improve system scalability, [115] tackles sub-channel assignment with multiple BSs as illustrated in Fig. 13c. The network structure is different from the scenario in which there is only one BS, when multiple BSs are introduced into the network. In this work, IRSs and users are coordinated by a BS and a server (such as a cellular station), but are not necessarily in the same coverage cell of the BS. When there are multiple BSs in different cells under assistance of multiple IRSs, the channel scheduling problem is formulated as a combinatorial optimization, which is also in the area of QUBO. And a QA based solution is proposed to solve the problem. However, it is a centralized scheduling method for multiple IRSs. In addition, there is no detailed discussion on QA implementation though the simulations validated its effectiveness.
@@
[114] and [112] consider network performance in a communication cell with multiple users, an IRS and one BS. The network is easily extended to a communication cell with multiple users, multiple IRSs, and multiple BSs in different cells. To reduce the network overhead such as hardware costs, it obtains the similar network performance with [115] by sharing an ISR of switching phase shift as illustrated in Fig. 13d, which is studied in [116]. Specifically, when a particular phase shift is set at a given time, the corresponding time resources and BSs are assigned to UEs existing in the reflection direction of the radio waves transmitted from each BS, which enables efficient communication. In the case of multiple operators, each operator performs such radio resource-allocation control in parallel, enabling communications with the IRS without coordination required between the operators. The process is formulated as QUBO and solved by classical QC algorithm. However, it can not achieve the optimal performance since multiple operators work independently. And thus, it is still viewed as a centralized approach. Other works related to IRS scheduling for channel assignment includes [113] that considers both power allocation and communication security. Its scenario is similar with [114, 112] in a communication cell. However, the solution for channel assignment of scheduling phase shift is independent with other optimization variables. A quantum-inspired bald eagle search algorithm is developed to solve the problem.
@@
In summary, QC-enabled algorithms offer a promising solution to enhance performance in channel assignment, as the overview of these studies presented in Table VI and QC details in Table VII.
@@ III-D1 Overview
User association in wireless networks refers to the process of determining which access point, BS, or sensor sink that a user device connects to for communication. This decision is crucial for optimizing network performance, ensuring efficient resource utilization and service quality [118]. The primary goal of user association is to provide seamless connectivity while maximizing network capacity and throughput. Various factors influence user association decision, including signal strength, available bandwidth, network load, user mobility, and QoS requirements. Note that network dynamics, particularly, sink (such as UAVs) mobility, are one of key features in user association, which is explored in numerous studies [119], such as in mobile sensor networks and UAV networks. Solutions encounter computational constraints like with a substantial number of devices and dynamic environment. The challenge can be addressed through the utilization of QC techniques. Thus, in this section, we primarily categorize our discussion according to two key criteria: QC algorithm category and mobility of BSs or sinks.
@@ III-D2 Applications of Conventional QC for User Association
TABLE VIII: Summary of QC Works on User Association in Wireless Networks
@@
The study examines basic wireless sensor network [120], consisting of sensors and a BS. It introduces a clustering method based on quantum evolution algorithm (QEA) to extend the network lifespan. However, it does not account for network dynamics, such as traffic patterns and device mobility. In [121], a comparable scenario to [120] is explored, with the introduction of the quantum-inspired gravitational search to enhance energy efficiency and prolong network lifespan. The study considers communication constraints and application-specific criteria, encompassing parameters such as mean relative deviation, spatial density error, and sensor-out-of-range error. Nonetheless, the paper lacks an in-depth analysis of the optimality of the proposed solution. Meanwhile, [122] presents an alternative quantum-inspired approach to the issue addressed in [120] and [121]. Named the quantum-inspired gravitational search algorithm, this solution aims to determine the operational states of sensors while preserving network connectivity. It introduces a weighted fitness function, integrating connectivity, energy, and application-specific parameters. However, accurately determining the appropriate weights for these factors poses a challenge.
Several additional quantum-inspired approaches have been proposed in the literature, including works such as [123, 124, 125, 126, 127, 128]. In the proposed quantum-inspired elite gray wolf approach [123], each individual in the population is represented by a set of quantum probability amplitudes corresponding to the states of all qubits. The initialization of the quantum probability amplitude matrix is performed by a logistic chaotic map, and the matrix is subsequently updated using quantum rotation gates based on the individual fitness value at each iteration. To enhance the search for optimal clustering, the method introduces an elite pool that stores the best-performing individuals and their corresponding quantum probability amplitude matrices from previous iterations. During the $g$-th iteration, top individuals are selected from the elite pool based on their fitness and are inherited to guide the population evolution in the subsequent $(g+1)$-th iteration. Additionally, a quantum NOT gate is applied to prevent the algorithm from converging prematurely to local optima.
These contributions introduce techniques such as quantum elite gray wolf optimization, quantum clone whale optimization, quantum artificial bee colony, QPSO, and QGA, all aimed at addressing the same problem outlined in [120]. Additionally, other studies explore link connectivity within QC algorithms. For example, [129] investigates online social networks like facebook, weibo, and twitter. The goal is to predict link connectivity, facilitating recommendations for registered users to connect with others.
@@
Subsequent studies consider more complex scenarios involving UAV networks and wireless sensor networks featuring mobile sinks. For example, [130] focuses on path planning of a mobile sink to enhance energy efficiency. In this scenario, the mobile sink starts its path from a designated location, traverses through the communication range of each sensor along the predetermined path, conducts a round of data collection, and returns to its initial starting point. They employ mathematical transformations to align the problem characteristics with those of a traveling salesman (TSP) problem. An optimal TSP path is obtained by using quantum tunneling effect and quantum circuit to achieve parallelism.
Regarding UAV networks, [131] considers a scenario with multiple UAVs but does not address data relay among UAVs. The research designs an objective function incorporating parameters like average distance, distance to neighboring UAVs, and the number of UAV neighbors to select cluster heads. While the proposal accounts for UAV mobility, it lacks mobility analysis within the model. Additionally, the paper explores UAV clustering application in image classification, but the model lacks a direct correlation between UAV clustering and image classification. In contrast, [132] tackles the challenge of UAV coalition formation for providing computation resources in a large-scale heterogeneous UAV network. The study operates in an unknown and dynamic environment, where UAVs dynamically acquire environmental knowledge. Leader UAVs assess follower UAV’ contributions in resource availability and share this information during coalition formation. QGA guides the inclusion of idle UAVs into coalitions, considering UAV mobility characterized by randomness and lacking correlation with ground points of interest (PoI). However, the study lacks an analysis of PoI impact on system performance, differing significantly from edge computing research, particularly in task offloading investigations [135].